When the radius of a circle is increased by two inches, its area is increased by 24 pi inches squared. What is the radius of the larger circle?
Answer by rothauserc(4718)

1) area(A) = pi * radius(r)^2
:
2) A +(24*pi) = pi * (r+2)^2
:
substitute for A in equation 2
:
(pi * r^2) +(24*pi) = pi * (r+2)^2
:
divide both sides of = by pi
:
r^2 + 24 = r^2 +4r +4
:
4r = 20
:
r = 5
:
*************************************************
The radius of the larger circle is 5+2 = 7 inches
*************************************************
: